# ClawMobile: Rethinking Smartphone-Native Agentic Systems

### v1

@@ Abstract.
$c$s: Computing methodologies Natural language processingccs: Human-centered computing Ubiquitous and mobile computingccs: Computer systems organization Embedded softwareccs: Software and its engineering Software performance
@@ 1. Introduction
  • We instantiate this architecture in ClawMobile and report preliminary results and observations from real-device deployments.
@@ 3.1. System Architecture
At the top of the stack, the Chat Channel serves as the entry point for high-level tasks. The orchestrator will interpret user input, generate task-level plans, and coordinate subsequent execution steps. The Agent Orchestrator does not directly manipulate device interfaces. Instead, it interacts with a set of Control Backends through explicit tool calls. These backends encapsulate concrete execution mechanisms, such as structured system commands (ADB), hardware-level APIs (Termux API (Termux-API, 2015)), or semantic UI interaction modules. Each backend exposes well-defined action primitives and returns bounded execution results. By routing execution through these backends, the orchestrator maintains a clear separation between reasoning and actuation.
@@ 3.4. Implementations
The Agent Orchestrator is built on top of OpenClaw’s agent framework, which provides task interpretation, tool invocation, and iterative planning capabilities. Control backends are implemented as modular components that expose structured action primitives. In our prototype, deterministic pathways are implemented via ADB-based system commands and hardware interfaces exposed through Termux API (Termux-API, 2015). For semantic UI interaction, we employ DroidRun (DroidRun, 2026) as the UI agent backend, leveraging its program-level action synthesis to support multi-step execution. These components collectively instantiate the heterogeneous control backends described in our runtime architecture. Each backend returns bounded, machine-readable results to the orchestrator, ensuring explicit execution semantics and avoiding implicit inference of device state. The runtime maintains a lightweight Memory component that stores execution preferences, reusable task patterns, and mobile-specific operational knowledge. This component influences backend selection and task decomposition but remains decoupled from low-level control logic. This knowledge is embedded into the context of model or stored in the agent’s retrieval path through OpenClaw’s memory mechanism.
@@ 5.2. Adaptability: Structuring Behavior in Heterogeneous App Ecosystems
Smartphone-native agents must operate across diverse and evolving application ecosystems. Adaptability therefore extends beyond reasoning accuracy to structural mechanisms that support reuse, specialization, and long-term coherence.
@@ 5.3. Stability: Reliability, Fault Tolerance, and Privacy
Reliability and Fault Tolerance. Mobile agents often encounter partial or silent failures, e.g., interrupted control channels, actions that do not take effect, or delayed state propagation, without explicit error signals. Reliable runtimes therefore need explicit progress verification, bounded execution scopes with well-defined outcomes, and principled recovery mechanisms (e.g., handling prompts, restoring focus, relaunching, or re-navigating) to re-establish control and resume safely. Elevating reliability to a first-class runtime objective, rather than an implementation detail, is essential for moving from prototypes to deployable agents.
Security and Privacy in Mobile Autonomy. Smartphone-native agents operate over sensitive user data and personal applications, so execution policies must incorporate privacy-aware state handling, bounded logging, and explicit permission mediation. Hybrid inference strategies further introduce trust boundaries between local and remote components, raising questions about what state is exported, under what guarantees, and how user control is preserved. Future work should formalize security and privacy models for mobile agent runtimes to ensure autonomy does not compromise data confidentiality or user intent.
@@ 6. Conclusion
By framing mobile agents as a runtime systems problem, this work shifts attention from isolated algorithmic advances toward integrated architectural design. Taken together, the three dimensions discussed—efficiency, adaptability, and stability—suggest that smartphone-native agentic systems demand a unified runtime perspective. Rather than optimizing individual components in isolation, future research must co-design reasoning, control, memory, and reliability under real-device constraints. We believe this perspective lays the groundwork for advancing mobile LLM autonomy beyond UI-centric automation toward principled, smartphone-native agent architectures.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Sixth European Workshop on Machine Learning and Systems ; April 27–30, 2026; Edinburgh, Scotland Ukbooktitle: Sixth European Workshop on Machine Learning and Systems (EuroMLSys ’26), April 27–30, 2026, Edinburgh, Scotland Ukdoi: 10.1145/3805621.3807655isbn: 979-8-4007-2605-7/2026/04ccs: Computing methodologies Natural language processingccs: Human-centered computing Ubiquitous and mobile computingccs: Computer systems organization Embedded softwareccs: Software and its engineering Software performance
@@ 1. Introduction
  • We instantiate this architecture in ClawMobile and report preliminary results and observations of real-device deployments.
@@ 3.1. System Architecture
At the top of the stack, the Chat Interface serves as the entry point for high-level tasks. The orchestrator will interpret user input, generate task-level plans, and coordinate subsequent execution steps. The Agent Orchestrator does not directly manipulate device interfaces. Instead, it interacts with a set of Control Backends through explicit tool calls. These backends encapsulate concrete execution mechanisms, such as structured system commands (ADB), hardware-level APIs (Termux API (Termux-API, 2015)), or semantic UI interaction modules. Each backend exposes well-defined action primitives and returns bounded execution results. By routing execution through these backends, the orchestrator maintains a clear separation between reasoning and actuation.
@@ 3.4. Implementations
The Agent Orchestrator is built on top of OpenClaw’s agent framework, which provides task interpretation, tool invocation, and iterative planning capabilities. Control Backends are implemented as modular components that expose structured action primitives. In our prototype, deterministic pathways are implemented via ADB-based system commands and hardware interfaces exposed through Termux API (Termux-API, 2015). For semantic UI interaction, we employ DroidRun (DroidRun, 2026) as the UI agent backend, leveraging its program-level action synthesis to support multi-step execution. These components collectively instantiate the heterogeneous control backends described in our runtime architecture. Each backend returns bounded, machine-readable results to the orchestrator, ensuring explicit execution semantics and avoiding implicit inference of device state. The runtime maintains a lightweight Memory component that stores execution preferences, reusable task patterns, and mobile-specific operational knowledge. This component influences backend selection and task decomposition but remains decoupled from low-level control logic. This knowledge is embedded into the context of the model or stored in the agent’s retrieval path through OpenClaw’s memory mechanism.
@@ 5.2. Adaptability: Structuring Behavior in Heterogeneous App Ecosystems
Smartphone-native agents must operate across diverse and evolving application ecosystems. Adaptability extends beyond reasoning accuracy to structural mechanisms that support reuse, specialization, and long-term coherence.
@@ 5.3. Stability: Reliability, Fault Tolerance, and Privacy
Reliability and Fault Tolerance. Mobile agents often encounter partial or silent failures, e.g., interrupted control channels, actions that do not take effect, or delayed state propagation, without explicit error signals. Reliable runtimes therefore need explicit progress verification, bounded execution scopes with well-defined outcomes, and principled recovery mechanisms (e.g., handling prompts, restoring focus, relaunching, or re-navigating) to re-establish control and resume safely. Elevating reliability to a first-class runtime objective, rather than an implementation detail, is essential to move from prototypes to deployable agents.
Security and Privacy in Mobile Autonomy. Smartphone-native agents operate over sensitive user data and personal applications, so execution policies must incorporate privacy-aware state handling, bounded logging, and explicit permission mediation. Hybrid inference strategies further introduce trust boundaries between local and remote components, raising questions about what state is exported, under what guarantees, and how user control is preserved. Future work should formalize security and privacy models for mobile agent runtimes to ensure that autonomy does not compromise data confidentiality or user intent.
@@ 6. Conclusion
By framing mobile agents as a runtime systems problem, this work shifts attention from isolated algorithmic advances toward integrated architectural design. The three dimensions discussed, efficiency, adaptability, and stability, suggest that smartphone-native agentic systems demand a unified runtime perspective. Rather than optimizing individual components in isolation, future research must co-design reasoning, control, memory, and reliability under real-device constraints. We believe this perspective lays the groundwork for advancing mobile LLM autonomy beyond UI-centric automation toward principled, smartphone-native agent architectures.